# spacing=nonfrench

### v1

@@ 5.4 The Equality $\mathcal{B}_{p,p}^{s}(\Omega\!\to\!\Gamma)=B_{p,p}^{s}(\Gamma)$
then it holds that $\mathcal{B}_{p,p}^{s}(\Omega\!\to\!\Gamma)=B_{p,p}^{s}(\Gamma)$ with comparable norms.
We recall that if $\delta$ stands for the infinimum of the set of $q$ such that $|\varphi^{\prime}|\in A_{q}$ then the conditions in the theorem are equivalent to
$\delta<q<\frac{\delta}{\delta-1}.$
In particular, $\delta<2$ so that $|\varphi^{\prime}|\in A_{2}$. That implies that $B_{p,p}^{s}(\Gamma)$ is stable by conjugation by Theorem 5.1. Together with $\text{Arg}\,\varphi^{\prime}\in L^{\infty}$, that also implies that $\mathcal{B}_{2,2}^{s}(\Omega\!\to\!\Gamma)=B_{2,2}^{s}(\Gamma)$, $0\leq s\leq 1$, with comparable norms by Theorem 5.5. These two results will be used in the proof of Theorem 5.7.
On the other hand, for the proof of Theorem 5.7 we also need to use the claim: $\mathcal{B}_{p,p}^{1/p}(\Omega\!\to\!\Gamma)=B_{p,p}^{1/p}(\Gamma)$ with comparable norms for $1<p<\infty$ provided that the curve $\Gamma$ is chord-arc. That was proved in our recent paper [45] using the conformal invariance of the space $\mathcal{B}_{p,p}^{1/p}(\Omega)$. We outline its proof here: suppose $f\in B_{p,p}^{1/p}(\Gamma)$ and its harmonic extension in $\Omega$ is still denoted by $f$ then the assertion follows from the following reasoning:
@@
where the implicit constants depend only on $\Omega$ and $p$. The above steps in order are due to $(1)$ $\lambda$ is a bi-Lipschitz homeomorphism of $\mathbb{T}$ onto $\Gamma$; $(2)$ $\lambda^{-1}\circ\varphi$ is a quasisymmetry; $(3)$ when $\Gamma=\mathbb{T}$ this is well-known as we mentioned before; $(4)$ the space $\mathcal{B}_{p,p}^{1/p}(\Omega)$ is conformal invariant.
Recall that $\mathcal{B}_{p,p}^{s}(\Omega)$ is conjugate invariant. We have seen, under the assumption of Theorem 5.7, that $B_{p,p}^{s}(\Gamma)$ is also conjugate invariant, in order to prove Theorem 5.7 we thus only need to prove the holomorphic case:
@@
$\mathcal{HB}_{p,p}^{s}(\Omega\!\to\!\Gamma)=\text{HB}_{p,p}^{s}(\Gamma)$ (5.11)
with comparable norms. The conclusion holds in particular for radial-Lipschitz domains.
@@
For any $f\in\mathcal{HB}_{p,p}^{s}(\Omega)$, $0<s<1$, $1<p<\infty$ we have seen from (5.2) that
@@
with comparable norms. Noting that any $g\in A^{p}_{(1-s)p-1}$ can be written in the form $V_{s,p}(f)^{\prime}$ by choosing $f(z)=\int_{0}^{z}g\circ\varphi^{-1}(u)(\varphi^{-1})^{\prime}(u)^{1+1/p-s}du$, we have that the operator $\mathcal{HB}_{p,p}^{s}(\Omega)\ni f\mapsto V_{s,p}(f)^{\prime}\in A^{p}_{(1-s)p-1}$ is a bounded isomorphism. This boils down to prove
@@
We have proved that for any $1<p<\infty$, $\mathcal{B}_{p,p}^{1/p}(\Omega\!\to\!\Gamma)=B_{p,p}^{1/p}(\Gamma)$ if $\Gamma$ is a chord-arc curve; and for any $0<s<1$, $\mathcal{B}_{2,2}^{s}(\Omega\!\to\!\Gamma)=B_{2,2}^{s}(\Gamma)$ if moreover $\Gamma$ satisfies that $|\varphi^{\prime}|\in A_{2}$ and $\text{Arg}\varphi^{\prime}\in L^{\infty}$, thus the equivalence (5.13) has already been known to hold for $s=1/p$, $1<p<\infty$, and also holds for $0<s<1$, $p=2$.
Recall that $(A^{p}_{p-2},A^{2}_{1-2s})$ is an interpolation pair and for $0\leq\theta\leq 1$,
@@
We have also seen that $(\text{HB}_{p,p}^{1/p}(\Gamma),\text{HB}_{2,2}^{s}(\Gamma))$ is an interpolation pair and
$\text{HB}_{p,p}^{1/p}(\Gamma),\text{HB}_{2,2}^{s}(\Gamma)]_{\theta}=\text{HB}_{q,q}^{\alpha}(\Gamma)$ (5.15)
@@
Some elementary arithmetic leads to that for any given $\alpha\in(0,1)$ and $q\in(1,\infty)$, there exist $p\in(1,\infty)$, $s\in(0,\,\min(1,5/2-2/p))$ and $\theta\in[0,1]$ such that (5.16) holds. Recall that in the proof of Theorem 5.5 we have shown that $\check{B}$, the set of polynomials on $\Gamma$, is dense in $\text{HB}_{2,2}^{s}(\Gamma)$. By using the similar argument, we can see that $\check{B}$ is also dense in $\text{HB}_{p,p}^{1/p}(\Gamma)$ and thus $\check{B}$ is a dense subspace of the intersection space $\text{HB}_{p,p}^{1/p}(\Gamma)\cap\text{HB}_{2,2}^{s}(\Gamma)$.
Define a family of linear mappings $(\widetilde{T}_{z};\;z=\theta+it\in S)$, $S$ standing for a vertical strip with $0\leq\theta\leq 1$ as before, on $\check{B}$ by
@@
We first estimate $|\widetilde{T}_{z}(f)(\zeta)|^{p}$ for each $z=\theta+it\in S$ and each $f\in\check{B}$. It follows from simple computation that
@@
There exists a constant $C_{1}$ such that $|f^{\prime}\circ\varphi|^{p}<C_{1}$ since the diameter of the domain $\Omega$ with 0 as an interior point is finite. Then,
@@
To show the linear mapping $\widetilde{T}_{z}$ sends $\check{B}$ to $A^{p}_{p-2}+A^{2}_{1-2s}$, we split into the following three cases:
@@
$\iint_{\mathbb{D}}|\varphi^{\prime}(z)|^{2+p\theta(\frac{1}{2}-s)}d(\varphi(z),\Gamma)^{p-2}dxdy<\infty.$
First of all, since $s>1/2$ and $p<2$ we have that
$1<2+p\theta(\frac{1}{2}-s)<2.$
@@
and the last integral is equal, by the change of variable $\zeta=\varphi(z)$, to
@@
A sufficient condition for the finiteness will thus be that
@@
Consequently, when $1<p<\infty,\,0<s<\min(1,5/2-2/p)$, the linear ma??pping $\widetilde{T}_{z}$ sends $\check{B}$ to $A^{p}_{p-2}+A^{2}_{1-2s}$, and for all $f\in\check{B}$, the mapping $\widetilde{T}_{(\cdot)}(f):S\to A^{p}_{p-2}+A^{2}_{1-2s}$ is bounded.
@@
By the continuity of parametrized integral, all functions $S\ni z\mapsto\widetilde{T}_{z}f\in A^{p}_{p-2}+A^{2}_{1-2s}$,   $\mathbb{R}\ni t\mapsto\widetilde{T}_{it}f\in A^{p}_{p-2}$, and $\mathbb{R}\ni t\mapsto\widetilde{T}_{1+it}f\in A^{2}_{1-2s}$ are continuous. Finally, for any $f\in\check{B}$, the proof of holomorphy of the mapping $\stackrel{{\scriptstyle\circ}}{{S}}\ni z\mapsto\widetilde{T}_{z}f\in A^{p}_{p-2}+A^{2}_{1-2s}$ is similar to that in the proof of Theorem 5.5. We omit the details here. This completes the proof of Theorem 5.8. ∎
5.5 Some Final Remarks
In this part, we let $0<s<1$ and $1<p<\infty$, and suppose $\Gamma$ satisfies the assumption of Theorem 5.7, namely $\Gamma$ is a bounded chord-arc curve with 0 in its interior domain $\Omega$, and $\Gamma$ is such that $\text{Arg}\,\varphi^{\prime}\in L^{\infty}$ and $|\varphi^{\prime}|\in A_{q}$ with $q=\min{(p,p^{\prime})}$. Here, $1/p+1/p^{\prime}=1$ and $\varphi_{i}^{\prime}$ is a Riemann mapping from the unit disk $\mathbb{D}$ onto $\Omega$ keeping 0 fixed. Recall that when $p=2$ we have shown that the radial-Lipschitz curve satisfies this assumption.
Since $h(\Gamma)=1$ one may see that $(h(\Gamma),p,s)$ is always in the admissible set $\mathcal{A}$. It follows from Theorem 6.2 and Theorem 5.7 (or Theorem 5.5 for the case of $p=2$) that in this case $\mathcal{B}_{p,p}^{s}(\Omega_{i}\!\to\!\Gamma)=\mathcal{B}_{p,p}^{s}(\Omega_{e}\!\to\!\Gamma)$ with comparable norms. Recall that we explored earlier that this equality also holds under a different assumption where there are some restrictions on $p$ and $s$ in Theorem 3.12.
If $p>2$, $s\geq 1-1/p$ we then have that
  • (i)
    $\mathcal{B}_{p,p}^{s}(\Gamma)=W^{1,p}(\omega,\mathbb{C})|_{\Gamma}$ where $\omega(z)=d(z,\Gamma)^{(1-s)p-1}$, i.e., the space $\mathcal{B}_{p,p}^{s}(\Gamma)$ is the trace space of $W^{1,p}(\omega,\mathbb{C})$ on $\Gamma$;  (by Theorem 3.16)
  • (ii)
    The space $C_{c}^{\infty}(\mathbb{C})|_{\Gamma}$ is dense in $\mathcal{B}_{p,p}^{s}(\Gamma)$, and thus Plemelj-Calderón property holds for any function in $\mathcal{B}_{p,p}^{s}(\Gamma)$.  (by Corollary 3.17 and Theorem 3.5)
  • (iii)
    for any $f\in\mathcal{B}_{p,p}^{s}(\Gamma)$, its extension to $\mathbb{C}$ by the harmonic extensions on both connected components $\Omega_{i,e}$ of $\overline{\mathbb{C}}\setminus\Gamma$ belongs to $W^{1,p}(\omega,\mathbb{C})$;  (by section 3.3.2)
  • (iv)
    Any function in $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$ satisfies the almost-Dirichlet principle.  (by section 3.2.3)
The above statements are claims: given $p>2$ and thus the class of curves $\Gamma$ for which claims are true depends on $p$ (i,e., $|\varphi^{\prime}|\in A_{p^{\prime}}$). If we reverse the point of view and start with a curve $\Gamma$ satisfying $\text{Arg}\,\varphi^{\prime}\in L^{\infty}$ and $|\varphi^{\prime}|\in A_{2}$ then the claims are true for $p\in(\delta,\frac{\delta}{\delta-1})$ where $\delta$ is the infinimum of the set of $p^{\prime}s$ such that $|\varphi^{\prime}|\in A_{p}$. By the theory of $A_{\infty}$ weights, $\delta<2$.

### v2

@@ 5.4 The equality $\mathcal{B}_{p,p}^{s}(\Omega)=B_{p,p}^{s}(\Gamma)$
then it holds that $\mathcal{B}_{p,p}^{s}(\Omega)=B_{p,p}^{s}(\Gamma)$ with comparable norms.
We recall that if $\delta=\inf\{q>1:\,|\varphi^{\prime}|\in A_{q}\}$ then the theorem is true for
$\delta<p<\frac{\delta}{\delta-1}.$
In particular, $\delta<2$ so that $|\varphi^{\prime}|\in A_{2}$. That implies that $B_{p,p}^{s}(\Gamma)$ is stable by conjugation by Theorem 5.1. Together with $\text{Arg}\,\varphi^{\prime}\in L^{\infty}$, this also implies that $\mathcal{B}_{2,2}^{s}(\Omega)=B_{2,2}^{s}(\Gamma)$, $0\leq s\leq 1$, with comparable norms by Theorem 5.5. These two results will be used in the proof of Theorem 5.7.
On the other hand, for the proof of Theorem 5.7 we also need to use the claim: $\mathcal{B}_{p,p}^{1/p}(\Omega)=B_{p,p}^{1/p}(\Gamma)$ with comparable norms for $1<p<\infty$ provided that the curve $\Gamma$ is chord-arc (see Theorem 1.1). For convenience we sketch its proof here: suppose that $f\in B_{p,p}^{1/p}(\Gamma)$ and its harmonic extension in $\Omega$ is still denoted by $f$ then the assertion follows from the following reasoning:
@@
where the implicit constants depend only on $\Omega$ and $p$. The above steps are in order due to $(1)$ $\lambda$ is a bi-Lipschitz homeomorphism of $\mathbb{T}$ onto $\Gamma$; $(2)$ $\lambda^{-1}\circ\varphi$ is a quasisymmetry; $(3)$ when $\Gamma=\mathbb{T}$ this is well-known, as we mentioned before; $(4)$ the space $\mathcal{B}_{p,p}^{1/p}(\Omega)$ is conformal invariant.
Recall that $\mathcal{B}_{p,p}^{s}(\Omega)$ is conjugate invariant. We have seen, under the assumption of Theorem 5.7, that $B_{p,p}^{s}(\Gamma)$ is also conjugate invariant (see Theorem 5.1), in order to prove Theorem 5.7 we thus only need to prove the following
@@
$\mathcal{HB}_{p,p}^{s}(\Omega)=\text{HB}_{p,p}^{s}(\Gamma)$ (5.11)
with comparable norms. The conclusion holds in particular for the radial-Lipschitz domains.
@@
For any $f\in\mathcal{HB}_{p,p}^{s}(\Omega)$, $0<s<1$, $1<p<\infty$, we have seen from (5.2) that
@@
with comparable norms. Noting that any $g\in A^{p}_{(1-s)p-1}$ can be written in the form $V_{s,p}(f)^{\prime}$ by choosing $f(z)=\int_{0}^{z}g\circ\varphi^{-1}(u)(\varphi^{-1})^{\prime}(u)^{1+1/p-s}du$, we find that the operator $\mathcal{HB}_{p,p}^{s}(\Omega)\ni f\mapsto V_{s,p}(f)^{\prime}\in A^{p}_{(1-s)p-1}$ is a bounded isomorphism. This boils down to proving the following:
@@
We have proved that for any $1<p<\infty$, $\mathcal{HB}_{p,p}^{1/p}(\Omega)=\text{HB}_{p,p}^{1/p}(\Omega)$ if $\Gamma$ is a chord-arc curve; and for any $0<s<1$, $\mathcal{HB}_{2,2}^{s}(\Omega)=\text{HB}_{2,2}^{s}(\Omega)$ if, moreover, $\Gamma$ satisfies $\text{Arg}\varphi^{\prime}\in L^{\infty}$, thus the equivalence (5.13) is already known for $s=1/p$, $1<p<\infty$, and also for $0<s<1$, $p=2$.
Recall that $(A^{p}_{p-2},A^{2}_{1-2s})$ is an interpolation pair, and for $0\leq\theta\leq 1$,
@@
We have also seen that $(\text{HB}_{p,p}^{1/p}(\Omega),\text{HB}_{2,2}^{s}(\Omega))$ is an interpolation pair and
$\text{HB}_{p,p}^{1/p}(\Omega),\text{HB}_{2,2}^{s}(\Omega)]_{\theta}=\text{HB}_{q,q}^{\alpha}(\Omega)$ (5.15)
@@
Some elementary arithmetic leads to the conclusion that for any given $\alpha\in(0,1)$ and $q\in(1,\infty)$, there exist $p\in(1,\infty)$, $s\in(0,\,\min(1,5/2-2/p))$, and $\theta\in[0,1]$ such that (5.16) holds. Recall that in the proof of Theorem 5.5 we have shown that $\check{B}$, the set of polynomials on $\Gamma$, is dense in $\text{HB}_{2,2}^{s}(\Omega)$. Using a similar argument, we can see that $\check{B}$ is also dense in $\text{HB}_{p,p}^{1/p}(\Omega)$ and thus $\check{B}$ is a dense subspace of the intersection space $\text{HB}_{p,p}^{1/p}(\Omega)\cap\text{HB}_{2,2}^{s}(\Omega)$.
Define a family of linear mappings $(\widetilde{T}_{z};\;z=\theta+it\in S)$, $S$ representing a vertical strip with $0\leq\theta\leq 1$ as before, on $\check{B}$ by
@@
We first estimate $|\widetilde{T}_{z}(f)(\zeta)|^{p}$ for each $z=\theta+it\in S$ and each $f\in\check{B}$. From a simple computation, it follows that
@@
There exists a constant $C_{1}$ such that $|f^{\prime}\circ\varphi|^{p}<C_{1}$ since the diameter of the domain $\Omega$, with 0 as an interior point, is finite. Then
@@
To show that the linear mapping $\widetilde{T}_{z}$ sends $\check{B}$ to $A^{p}_{p-2}+A^{2}_{1-2s}$, we split into the following three cases:
@@
$\iint_{\mathbb{D}}|\varphi^{\prime}(z)|^{2+p\theta\left(\frac{1}{2}-s\right)}d(\varphi(z),\Gamma)^{p-2}dxdy<\infty.$
First of all, since $s>1/2$ and $p<2$ we have
$1<2+p\theta\left(\frac{1}{2}-s\right)<2.$
@@
and the last integral is equal, by the change of variables $\zeta=\varphi(z)$, to
@@
Thus, a sufficient condition for finiteness will be that
@@
Consequently, when $1<p<\infty,\,0<s<\min(1,5/2-2/p)$, the linear mapping $\widetilde{T}_{z}$ sends $\check{B}$ to $A^{p}_{p-2}+A^{2}_{1-2s}$, and for all $f\in\check{B}$, the mapping $\widetilde{T}_{(\cdot)}(f):S\to A^{p}_{p-2}+A^{2}_{1-2s}$ is bounded.
@@
By the continuity of the parametrized integral, all functions $S\ni z\mapsto\widetilde{T}_{z}f\in A^{p}_{p-2}+A^{2}_{1-2s}$,   $\mathbb{R}\ni t\mapsto\widetilde{T}_{it}f\in A^{p}_{p-2}$, and $\mathbb{R}\ni t\mapsto\widetilde{T}_{1+it}f\in A^{2}_{1-2s}$ are continuous. Finally, for any $f\in\check{B}$, the proof of holomorphy of the mapping $\stackrel{{\scriptstyle\circ}}{{S}}\ni z\mapsto\widetilde{T}_{z}f\in A^{p}_{p-2}+A^{2}_{1-2s}$ is similar to that in the proof of Theorem 5.5. We omit the details here. This completes the proof of Theorem 5.8. ∎
Remark 5.9.
Recall that in Theorems 5.5 and 5.7 we only focused on the case of bounded domains. The case of unbounded domains can actually be transformed to the bounded case; see Appendix for details, and similar assertions also hold for the unbounded case.
Let $\Gamma$ be a radial Lipschitz curve and $\Omega_{i,e}$ be the inner domain and the outer domain of $\Gamma$. Let $\varphi_{i,e}$ be a Riemann map for $\Omega_{i,e}$. By Helson-Szegö theorem, we see $|\varphi_{i,e}^{\prime}|\in A_{p}$, $p<2$ on $\mathbb{T}$. Set
$p_{0}^{i,e}=\inf\{p>1:\;|\varphi_{i,e}^{\prime}|\in A_{p}\}\;\;\mathrm{and}\;\;p_{0}=\max(p_{0}^{i},p_{0}^{e}).$
We end this subsection with an estimate of $p_{0}$ in terms of the geometry of the curve. Notice that the two functions
$\text{Arg}\left(\frac{z\varphi^{\prime}_{i,e}(z)}{\varphi_{i,e}(z)}\right)$
have the same $L^{\infty}$-norm that is less than $\pi/2$. We define $M$ as being the $\tan$ of this norm and call this number the Lipschitz-norm of $\Gamma$. We now use the same argument as that below Theorem 5.5:
$\log|\varphi_{i,e}^{\prime}(z)|^{p}=\log\left|\frac{\varphi_{i,e}(z)}{z}\right|^{p}+\log\left|\frac{z\varphi_{i,e}^{\prime}(z)}{\varphi_{i,e}(z)}\right|=\log\left|\frac{\varphi_{i,e}(z)}{z}\right|^{p}-H\left(p\text{Arg}\frac{z\varphi_{i,e}^{\prime}(z)}{\varphi_{i,e}(z)}\right),$
we thus conclude by Helson-Szegö theorem that $|\varphi_{i,e}^{\prime}|^{p}$ is an $A_{2}$-weight on the unit circle $\mathbb{T}$ if $p<\frac{\pi}{2\arctan M}$, and this holds in particular for $p=1$. On the other hand, let $\omega$ be an $A_{2}$-weight and $\tilde{p}$ the supremum of the set of $p$ such that $\omega^{p}\in A_{2}$. It is not difficult to see that $\delta(\omega)(:=\inf\{q>1:\,\omega\in A_{q}\})\leq 1+\frac{1}{\tilde{p}}$. Then we get an estimate value of $p_{0}$:
$1+\frac{1}{\frac{\pi}{2\arctan M}}=1+\frac{2\arctan M}{\pi}.$
We may then state the following corollary of Theorem 5.7.
Corollary 5.10.
If $\Gamma$ is a $M$-radial-Lipschitz curve and $p\in\left(1+\frac{2\arctan M}{\pi},1+\frac{\pi}{2\arctan M}\right)$, $s\in(0,1)$ then
Notice that this interval of $p$-value converges to $(1,\infty)$ as $M\to 0$ and to a singleton $\{2\}$ as $M\to\infty$.
5.5 Almost-Dirichlet principle revisited.
In Section 3 we have proven Theorem 3.5 which states that given a quasicircle $\Gamma,\,p>2,\,$ $\frac{h(\Gamma)}{2}-\frac{h(\Gamma)-1}{p}<s<1-\frac{h(\Gamma)-1}{p}$ (notice that when $\Gamma$ is chord-arc $h(\Gamma)=1$ and thus $1/2<s<1$) then if
$\mathcal{B}_{p,p}^{s}(\Omega_{i}\to\Gamma)=\mathcal{B}_{p,p}^{s}(\Omega_{e}\to\Gamma)$
we then have that this space, which we have denoted by $\mathcal{B}_{p,p}^{s}(\Gamma)$, is equal to $W^{1,p}(\mathbb{C},\omega)|_{\Gamma}$. Moreover, we have shown that the inclusion $W^{1,p}(\mathbb{C},\omega)|_{\Gamma}\subset\mathcal{B}_{p,p}^{s}(\Gamma)$ implies that the spaces $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$ obey the almost-Dirichlet principle, which means that there exists $C>0$ such that for any continuous function $F$ on $\overline{\Omega}$ which is a $C^{1}$-function in $\Omega$, if we call $u$ the harmonic extension of $F|_{\Gamma}$ in $\Omega$ then
$\|u\|_{\mathcal{B}_{p,p}^{s}(\Omega)}^{p}\leq C\iint_{\Omega}|\nabla F(z)|^{p}d(z,\Gamma)^{(1-s)p-1}dxdy.$
As before, let $\varphi_{i,e}$ be a Riemann map for $\Omega_{i,e}$, and let
$p_{0}^{i,e}=\inf\{p>1:\;|\varphi_{i,e}^{\prime}|\in A_{p}\}\;\;\mathrm{and}\;\;p_{0}=\max(p_{0}^{i},p_{0}^{e}).$
Theorem 5.11.
If $\Gamma$ is a radial-Lipschitz curve then $\mathcal{B}_{p,p}^{s}(\Omega_{i})$ has the almost-Dirichlet property if $2<p<\frac{p_{0}}{p_{0}-1}$ and $s>1/2$.
Proof.
It follows immediately from Theorems 3.5, 5.1, 5.7. ∎
Using the estimate value of $p_{0}$ in the last subsection, we have the following corollary.
Corollary 5.12.
If $\Gamma$ is a radial-Lipschitz curve with norm $M>0$ then $\mathcal{B}_{p,p}^{s}(\Omega_{i})$ has the almost-Dirichlet property if
$2<p<1+\frac{\pi}{2\arctan M},\quad\frac{1}{2}<s<1.$
This interval of the $p$-value converges to $(2,\infty)$ as $M\to 0$ and to a singleton $\{2\}$ as $M\to\infty$.